OpenStudy (anonymous):

For what two values is this expression undefined? Type your answer as two integers separated by a comma, like this: -4, 12 x^2+x−2/x^2+9x−10

OpenStudy (mertsj):

The two values that cause the denominator to be 0.

OpenStudy (anonymous):

so -2, -10

OpenStudy (mertsj):

no

OpenStudy (mertsj):

Solve this equation: \[x^2+9x-10=0\]

OpenStudy (anonymous):

i am not sure. really am not good with math.

OpenStudy (mertsj):

Factor the left side.

OpenStudy (anonymous):

im not sure, i cant seem to find the numbers in the factor

OpenStudy (mertsj):

Think of two numbers: If you multiply them you get -10 What could the numbers possibly be?

OpenStudy (anonymous):

-5,2 and -2,5

OpenStudy (mertsj):

Are there any other possibilities?

OpenStudy (anonymous):

-10,1 and -1,10

OpenStudy (mertsj):

So you have 4 possibilities. Take each pair of numbers and add them together. Is one answer 9?

OpenStudy (anonymous):

yes the -1 and 10

OpenStudy (mertsj):

typo. the factors are x-1 and x+10

OpenStudy (mertsj):

In other words. x^2+9x-10=(x-1)(x+10)

OpenStudy (mertsj):

So if you set the denominator equal to 0 you have: \[x^2+9x-10=0\] \[(x-1)(x+10)=0\]

OpenStudy (mertsj):

If the product of factors is 0, then at least one of the factors has to be 0. Set each factor equal to 0 and solve the two equations.

OpenStudy (anonymous):

\[x-1=0\] \[x=1\] \[0\]

OpenStudy (anonymous):

right

OpenStudy (mertsj):

Yes. 1 is one of the numbers that will cause the denominator to be 0. Find the other one.

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

this is what I got from solving the factor

OpenStudy (mertsj):

Very good.

OpenStudy (anonymous):

so thats the answer

OpenStudy (mertsj):

There are two answers. They are 1 and -10